OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22 A)4(-8x - 9) - 3y = -22 B)4x - 3(-8x - 9) = -22 C)4(8x + 9) - 3y = -22 D)4x - 3(8x + 9) = -22

OpenStudy (unklerhaukus):

ok if your going to use the substitution method , first you will have to solve the first equation for y , can you do that?

OpenStudy (anonymous):

y= 8x+9 right?

OpenStudy (unklerhaukus):

right

OpenStudy (unklerhaukus):

now subsite this into the second equation , what do you get

OpenStudy (anonymous):

D) 4x-3(8x +9) = -22

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (unklerhaukus):

now to continue to solve the system we can solve this new equation of x

OpenStudy (unklerhaukus):

when we have done that we substitute this back into the first equation and solve for y

OpenStudy (nirmalnema):

x=1/4.....y=11

OpenStudy (anonymous):

alright. Thank you!!

OpenStudy (nirmalnema):

ur welcome.....

OpenStudy (unklerhaukus):

almost @nirmalnema

OpenStudy (unklerhaukus):

but you'll find x=-1/4